# CloudCast – Total Cloud Cover Nowcasting with Machine Learning

### v1

@@ Abstract
Cloud cover plays a critical role in weather prediction and impacts several sectors, including agriculture, solar power generation, and aviation. Despite advancements in numerical weather prediction (NWP) models, forecasting total cloud cover remains challenging due to the small-scale nature of cloud formation processes. In this study, we introduce CloudCast, a convolutional neural network (CNN) based on the U-Net architecture, designed to predict total cloud cover (TCC) up to five hours ahead. Trained on five years of satellite data, CloudCast significantly outperforms traditional NWP models and optical flow methods. Compared to a reference NWP model, CloudCast achieves a 24% lower mean absolute error and reduces multi-category prediction errors by 46%. The model demonstrates strong performance, particularly in capturing the large-scale structure of cloud cover in the first few forecast hours, though later predictions are subject to blurring and underestimation of cloud formation. An ablation study identified the optimal input features and loss functions, with MAE-based models performing the best. CloudCast has been integrated into the Finnish Meteorological Institute’s operational nowcasting system, where it improves cloud cover forecasts used by public and private sector clients. While CloudCast is limited by a relatively short skillful lead time of about three hours, future work aims to extend this through more complex network architectures and higher-resolution data. CloudCast code is available at https://github.com/fmidev/cloudcast.
@@ 1 Introduction
Clouds play a vital role in Earth’s atmospheric energy balance by regulating temperature, distributing moisture, and influencing weather patterns (Quante, 2004). Their effects extend across various sectors, including agriculture, solar power generation, aviation, and everyday life, where cloud cover forecasts help people plan activities (Lazo et al., 2009). However, predicting cloud cover accurately remains a challenge for numerical weather prediction (NWP) models, due to the small-scale nature of cloud formation (Tiedtke, 1993), the spatial and temporal variability of observational data (Dai et al., 2006), and the complexities of assimilating remote-sensing data into NWP models (Bauer et al., 2011). Nowcasting, defined as short-term forecasting of up to 6 hours ahead (Wang et al., 2017), relies on on frequently updated, high resolution observational data, which is typically extrapolated into short range forecasts using methods such as optical flow and neural networks. While precipitation nowcasting using radar information is well-established and traditional methods provide accurate forecasts up to three hours, neural networks have extended this horizon to 12 hours (Sønderby et al., 2020). However, cloud cover nowcasting is less developed because of the lack of high-quality observational data that directly describe cloud fraction. Cloud fraction cannot be easily derived from raw satellite measurements.
@@
Finnish Meteorological Institute (FMI) provides an hourly-updated, 10-day forecast used by other government agencies, private sector and public. The forecast is a combination of direct model output from numerical weather prediction (NWP) models, statistically calibrated data and observation extrapolation data. CloudCast is integrated to FMI operational weather production and hourly forecasts are available to all FMI users.
The rest of the paper is organized as follows: Section 2 describes the curation of training and validation data, the selection of the neural network model, ablation study, and the model training details. Section 3 presents the results from the ablation study and the verification results for CloudCast, including case studies. Finally, section 4 concludes the study.
1.1 Related work
Optical flow methods use sequences of images, such as precipitation maps, to estimate the movement of the field. They typically operate in two phases: first, consecutive frames of cloud cover are analyzed to generate an atmospheric motion vector (AMV) field, which describes the past movement of weather phenomena. This AMV field is applied to the current frame to extrapolate future states. The process is repeated, using each new forecast as input, but without recalculating the AMV field until the desired forecast length is achieved. Since these methods only advect the existing field and do not account for the growth or decay of clouds, their prediction horizon is limited to a few hours. Operational forecasts often combine optical flow methods with numerical weather prediction (NWP) models to produce seamless forecasts. Typically, the first one or two hours rely on optical flow methods. As the forecast quality from optical flow begins to decline, NWP data is used to continue the forecast (Blahak and Keller, 2021; Roberts et al., 2023).
In machine learning, convolutional neural networks (CNNs) use a hierarchical structure where information content is increasingly abstracted as the data is passed through the network. Typical models are autoencoders that contain a contracting path that captures high-level features and an expanding path that reconstructs the output. CNNs have are also becoming increasingly popular for sequence-to-sequence nowcasting tasks, such as predicting precipitation and cloud cover (Moskolaï et al., 2021). Various architectures have been successful in different studies, most commonly using either explicit-memory models (LSTM, GRU) or architectures where temporal relations are described implicitly, such as U-Net. Feng et al. (2022) developed a shallow CNN to predict solar irradiance from sky imager data. Berthomier et al. (2020) trained several different CNNs to predict a binary cloud mask, outperforming NWP forecasts. CNNs have also been combined with physics-informed modules (Montassir et al., 2023) as well as optical flow methods (Ritvanen et al., 2023). Explicit-memory networks have been exploited extensively in spatio-temporal prediction tasks using typically either LSTM (Nielsen et al. (2021), Shi et al. (2015), Kumar et al. (2020)), Xia et al. (2024)) or GRU (Kellerhals et al. (2022), Wang et al. (2023)) based memory.
@@ 2 Methods
This section formulates our ML approach by first discussing the dataset used for model training and selection (see Section 2-a). Then the proposed neural network model is described in Section 2-b. Finally, we discuss the loss functions and hyper-parameters used for model training in Section 2-c.
@@ 2.1 Data
Spinning Enhanced Visible and Infrared Imager (SEVIRI), an instrument on board Meteosat-10, provides data in 12 spectral bands (Aminou et al., 1997). This data, along with any required auxiliary data, is processed with NWCSAF software suite for geostationary satellites (NWC/GEO) (Kerdraon and Fontaine, 2021). As a result a direct estimate of the total cloud cover, a parameter called effective cloudiness, is produced. Effective cloudiness is a part of the cloud top temperature and height (CTTH) family and is defined as the cloud fraction multiplied by cloud emissivity in the 10.8µm window channel (Kerdraon and Fontaine, 2021). It has a value of 1 for thick clouds, 0 for cloud-free areas and a value between 1 and 0 for semi-transparent clouds. The parameter has been empirically determined to semantically match quite closely the TCC observed and forecasted (by NWP). Effective cloudiness is produced in real-time in 15-minute intervals as a part of the FMI operational systems supporting duty forecasters.
Effective cloudiness data from November 2018 to October 2023 was used for training, a total of 1795 days and 166574 individual images. For evaluating the model, data from November 2023, January, April and June 2024 was used, totalling to 121 days and 11598 images and covering all seasons. FMI archive for effective cloudiness data extends over three different NWC/GEO versions: v2016, v2018 and v2021; no significant change in the meteorological quality of the data between versions was found. There were gaps in the time series, due to issues either at EUMETSAT or FMI production systems, but they were not addressed in any way. The FMI data archive is quality-controlled, and aside from gaps in the time series, no traces of data corruption or other anomalous values were discovered. The histogram of the cloud cover training data is shown in Figure 2. The cloud cover data has strong bimodal distribution, where more than 60% of the data is either 0 or 1.
@@
NWCSAF software outputs the full disk data in Netcdf format in a geostationary satellite specific spatial projection. The data was reprojected to lambert conformal conic projection matching the domain of interest, and converted to grib file format. This geographical domain is located between 50°N and 75°N, 20°W and 45°E, as seen in Figure 1. The horizontal resolution of the data after reprojection was 4 km. In high latitudes the effective resolution is worse than this as the satellites view is less orthogonal with relation to the surface of the Earth. A parallax-corrected version of the data was used to reduce the distortion in the data. All missing values were replaced with zero and the data was scaled between 0 and 1.
The method developed in the study was benchmarked against the conventional methods. The first benchmark data is the output from the control member from MetCoOp Ensemble Prediction System (MEPS). MEPS is using Harmonie-AROME cycle 43h2.2 (Bengtsson et al., 2017), a non-hydrostatic, convection-permitting numerical weather prediction system run in cooperation by the meteorological institutes of Estonia, Finland, Norway and Sweden (Müller et al., 2017). MEPS forecast is produced every three hours and it provides an hourly forecast up to 66 hours, with a horizontal resolution of 2.5km. The cloud fraction is derived for each of the 65 vertical levels of the model using a statistical cloud scheme, and the total fraction is computed assuming a random overlap between successive layers. For verification the analysis field and the first five hours of the forecast were used. The second benchmark data was EXIM, an optical flow prediction system which is a part of the NWCSAF suite of products and (in FMI’s configuration) provides a 60-minute forecast in 15-minute prediction intervals (ZAMG, 2021). FMI archives contain EXIM data only for April and June 2024. The third benchmark data was Eulerian persistence: using the current cloudiness data as a (static) prediction field for all future predictions, which especially for the first intra-hour forecasts can be hard to beat.
@@ 2.2 Neural network
The CloudCast architecture was inspired by the U-Net model originally developed for image segmentation (Ronneberger et al., 2015). U-Net is an encoder-decoder convolutional neural network named after its symmetric U-shaped structure, which features a contracting path to capture context and an expanding path for localization. It employs skip connections, allowing input data to be directly transferred from the encoder to the decoder. Although initially developed for medical imaging, U-Net has proven to be a simple yet effective network, successfully applied to various domains. The advantages of U-Net over alternative convolutional networks include: 1) fast training due to its end-to-end structure and context-based learning, 2) skip connections that enable the reuse of low-level features, improving accuracy without requiring large datasets, and 3) the ability to capture both global context (via the encoder) and fine pixel-level details (via the decoder) (Siddique et al., 2021). The network was implemented for CloudCast using the TensorFlow framework (Abadi et al., 2016).
@@
The network architecture is illustrated in Figure 3. The input sequence of effective cloudiness images, lead time and date and time information is gradually contracted as it passes through the convolutional and max-pooling layers, allowing the encoder to extract increasingly abstract and high-level features. The backbone of the network is a convolutional layer, where the image size is reduced to 32x32 pixels, while the number of channels is increased to 1024. Conv2DTranspose layers are used to upsample the dense latent representation into a prediction. While input data is passed along the encoder it is also concatenated directly into the decoder path to ensure that high-level features are preserved in the output. The final layer is a sigmoid activation function, which constrains the output values between 0 and 1. The output is a pixel-wise prediction of cloud cover for the specified lead time. The model consists of 31 million trainable parameters.
The approach of Sønderby et al. (2020) was adopted, where the forecast lead time is concatenated with the input data sequence. This method allows each lead time to be forecasted independently, enabling parallelization of the entire process and ensuring that each prediction is conditioned solely on ground truth data, reducing aleatoric uncertainty. However, using lead time as an additional input feature has drawbacks: the maximum lead time becomes an input parameter that must be determined prior to training the model, and the model’s ability to predict further into the future is constrained, as it does not observe the evolution of the weather system. While an autoregressive model can theoretically generate infinitely long predictions, the temporal range for this model is intentionally limited. This is acceptable as the focus of this study is exclusively on the nowcasting range of up to five hours.
@@ 2.3 Training details
An ablation study was conducted to determine the optimal loss function and input features for the model. The study tested different numbers of sequential input images—1, 2, 4, and 6—representing temporal histories of 0, 15, 45, and 75 minutes, respectively. Additionally, various static datasets were incorporated, including topography, ground type, date and time, and sun elevation angle. The strong bimodal distribution of the cloud cover data made selecting the right loss function particularly crucial. Standard loss functions, such as mean absolute error and mean squared error, may not perform well, as they tend to minimize validation error by consistently predicting 50% cloud cover. To assess the effectiveness of the loss function, six different options were tested in the ablation study.
@@
Mean absolute error (MAE) and mean squared error (MSE), also known as L1 and L2 losses, are widely used in machine learning and serve as strong baselines for comparison with other loss functions. Binary cross-entropy (BCE) measures the difference between predicted and actual probability distributions and is typically used in binary classification tasks to predict the probability that an input belongs to one of two categories. In the case of CloudCast, the prediction is a two-dimensional array of cloud fraction values between 0 and 1. Notably, the distribution of these values is heavily concentrated at 0 and 1 (see Figure 2). Because the cloud fraction values predominantly resemble a binary distribution, BCE can be used as a loss function, despite the non-categorical nature of the cloud cover data. Jungersen et al. (2022) found in their study BCE+MAE, a combination of BCE and MAE, to outperform plaine BCE. Log-cosh falls between MAE and MSE; it is robust against outliers like MAE but penalizes smaller errors more aggressively. Lastly, the Structural Similarity Index Measure (SSIM) assesses the similarity between two images based on structure, brightness, and contrast. It computes scores for multiple patches using a sliding window mechanism and aggregates these to provide a single global value for the image pair (Wang et al., 2004). As a loss function, SSIM assesses the similarity between images from a human perspective, considering larger-scale structures in the data rather than just pixel-wise errors like MAE.
@@
The models trained in the ablation study are summarised in Table 1. Date and time were converted into cyclic variables using sine and cosine transformation. The terrain type was consolidated from 30 different categorical values into 6 major types: forest, urban, bare, permanent snow, water, agriculture, and other. The sun elevation angle was precomputed for each grid point and each hour of the year using an internal formula from FMI. The models were trained in a horizontal resolution of 20km.
Table 1: Models included in the ablation study. A total of 13 different models were trained.
@@
The best model was selected based on the results of the ablation study and retrained using a 5km horizontal resolution (corresponding to 512x512 image size) over the same time period. The dataset size was 166GB. All the model hyperparameters were kept same, except batch size, which was reduced to 4. The training time for this model was 6 days.

### v2

@@ Abstract
Cloud cover plays a critical role in weather prediction and impacts several sectors, including agriculture, solar power generation, and aviation. Despite advancements in numerical weather prediction (NWP) models, forecasting total cloud cover remains challenging due to the small-scale nature of cloud formation processes. In this study, we introduce CloudCast, a convolutional neural network (CNN) based on the U-Net architecture, designed to predict total cloud cover (TCC) up to five hours ahead. Trained on five years of satellite data, CloudCast significantly outperforms traditional NWP models and optical flow methods. Compared to a reference NWP model, CloudCast achieves a 24% lower mean absolute error and reduces multi-category prediction errors by 46%. The model demonstrates strong performance, particularly in capturing the large-scale structure of cloud cover in the first few forecast hours, though later predictions are subject to blurring and underestimation of cloud formation. Model selection was performed using sensitivity experiments which identified the optimal input features and loss functions, with MAE-based models performing the best. CloudCast has been integrated into the Finnish Meteorological Institute’s operational nowcasting system, where it improves cloud cover forecasts used by public and private sector clients. While CloudCast is limited by a relatively short skillful lead time of about three hours, future work aims to extend this through more complex network architectures and higher-resolution data. CloudCast code is available at https://github.com/fmidev/cloudcast.
@@ 1 Introduction
Clouds play a vital role in Earth’s atmospheric energy balance by regulating temperature, distributing moisture, and influencing weather patterns (Quante, 2004). Their effects extend across various sectors, including agriculture, solar power generation, aviation, and everyday life, where cloud cover forecasts help people plan activities (Lazo et al., 2009). However, predicting cloud cover accurately remains a challenge for numerical weather prediction (NWP) models, due to the small-scale nature of cloud formation (Tiedtke, 1993), the spatial and temporal variability of observational data (Dai et al., 2006), and the complexities of assimilating remote-sensing data into NWP models (Bauer et al., 2011). Nowcasting, defined as short-term forecasting of up to 6 hours ahead (Wang et al., 2017), relies on frequently updated, high resolution observational data, which is typically extrapolated into short range forecasts using methods such as optical flow and neural networks. While precipitation nowcasting using radar information is well-established and traditional methods provide accurate forecasts up to three hours, neural networks have extended this horizon to 12 hours (Sønderby et al., 2020). However, cloud cover nowcasting is less developed because of the lack of high-quality observational data that directly describe cloud fraction. Cloud fraction cannot be easily derived from raw satellite measurements.
Optical flow methods use sequences of images, such as precipitation maps, to estimate the movement of the field. They typically operate in two phases: first, consecutive frames of cloud cover are analyzed to generate an atmospheric motion vector (AMV) field, which describes the past movement of weather phenomena. This AMV field is applied to the current frame to extrapolate future states. The process is repeated, using each new forecast as input, but without recalculating the AMV field until the desired forecast length is achieved. Since these methods only advect the existing field and do not account for the growth or decay of clouds, their prediction horizon is limited to a few hours. Operational forecasts often combine optical flow methods with numerical weather prediction (NWP) models to produce seamless forecasts. Typically, the first one or two hours rely on optical flow methods. As the forecast quality from optical flow begins to decline, NWP data is used to continue the forecast (Blahak and Keller, 2021; Roberts et al., 2023). Machine learning methods can learn complex, nonlinear cloud dynamics, enabling them to capture both advection and formation/dissipation processes, whereas optical flow primarily models motion based on pixel displacement. Additionally, ML models can leverage historical patterns and multi-source data (e.g., satellite, reanalysis) to improve forecasts, while optical flow relies solely on recent frame sequences, limiting its ability to predict cloud evolution beyond short lead times.
In machine learning, convolutional neural networks (CNNs) use a hierarchical structure where information content is increasingly abstracted as the data is passed through the network. Typical models are autoencoders that contain a contracting path that captures high-level features and an expanding path that reconstructs the output. CNNs are also becoming increasingly popular for sequence-to-sequence nowcasting tasks, such as predicting precipitation and cloud cover (Moskolaï et al., 2021). Various architectures have been successful in different studies, most commonly using either explicit-memory models (LSTM, GRU) or architectures where temporal relations are described implicitly, such as U-Net. Feng et al. (2022) developed a shallow CNN to predict solar irradiance from sky imager data. Berthomier et al. (2020) trained several different CNNs, including a U-Net to predict a binary cloud mask, outperforming NWP forecasts. Similarly Lagerquist et al. (2021) used a U-Net to predict convection using multipectral images from Himawari-8, and Fan et al. (2024) used a ResNet (residual neural network) to predict convection initialization from GOES-16 satellite observations. CNNs have also been combined with physics-informed modules (Montassir et al., 2023) as well as optical flow methods (Ritvanen et al., 2023). In time series prediction, recurrent neural networks (RNN) have been utilized widely. These networks are designed to learn temporal dependencies in the data, with network layers that can model the temporal dimension present in the data. For spatio-temporal prediction tasks typically either LSTM (Nielsen et al., 2021; Shi et al., 2015; Kumar et al., 2020; Xia et al., 2024) or GRU (Kellerhals et al., 2022; Wang et al., 2023; Leinonen et al., 2022) based networks are used. The inclusion of memory layers in RNN-based networks increases their complexity and makes them more difficult to train compared to similar architectures that do not incorporate memory.
@@
Finnish Meteorological Institute (FMI) provides an hourly-updated, 10-day forecast used by other government agencies, private sector and public. The forecast is a combination of direct model output from numerical weather prediction (NWP) models, statistically calibrated data and observation extrapolation data. CloudCast is integrated to FMI’s operational weather production and hourly forecasts are available to all FMI users.
The rest of the paper is organized as follows: Section 2 describes the curation of training and validation data, the selection of the neural network model, the sensitivity experiments, and the model training details. Section 3 presents the results from the sensitivity experiments and the verification results for CloudCast, including case studies. Finally, section 4 concludes the study and outlines directions for future research.
@@ 2 Methods
This section formulates our ML approach by first discussing the dataset used for model training and selection (see Section 2a). Then the proposed neural network model is described in Section 2b. Finally, we discuss the loss functions and hyper-parameters used for model training in Section 2c.
@@ 2.1 Data
Spinning Enhanced Visible and Infrared Imager (SEVIRI), an instrument onboard Meteosat-10, provides data in 12 spectral bands (Aminou et al., 1997). This data, along with any required auxiliary data, is processed with NWCSAF software suite for geostationary satellites (NWC/GEO) (Kerdraon and Fontaine, 2021). As a result a direct estimate of the total cloud cover, a parameter called effective cloudiness, is produced. Effective cloudiness is a part of the cloud top temperature and height (CTTH) family and is defined as the cloud fraction multiplied by cloud emissivity in the 10.8µm window channel (Kerdraon and Fontaine, 2021). It has a value of 1 for thick clouds, 0 for cloud-free areas and a value between 1 and 0 for semi-transparent clouds. The parameter has been empirically determined to closely correspond to the TCC observed and forecasted (by NWP). Effective cloudiness is produced in real-time in 15-minute intervals as a part of the FMI operational systems supporting duty forecasters.
Effective cloudiness data from November 2018 to October 2023 was used for training, a total of 1795 days and 166574 individual images. For evaluating the model, data from November 2023, January, April and June 2024 was used, totalling to 121 days and 11598 images and covering all seasons. FMI archive for effective cloudiness data extends over three different NWC/GEO versions: v2016, v2018 and v2021; no significant change in the meteorological quality of the data between versions was found. There were gaps in the time series, due to issues either at EUMETSAT or FMI production systems, but they were not addressed in any way. The FMI data archive is quality-controlled, and aside from gaps in the time series, no traces of data corruption or other anomalous values were discovered. The histogram of the cloud cover training data is shown in Figure 2. The cloud cover data have a strong bimodal distribution, where more than 60% of the data is either zero or one.
@@
NWCSAF software outputs the full disk data in NetCDF format in a geostationary satellite specific spatial projection. The data was reprojected to lambert conformal conic projection matching the domain of interest, and converted to grib file format. This geographical domain is located between 50°N and 75°N, 20°W and 45°E, as seen in Figure 1. The horizontal resolution of the data after reprojection was 4 km. In high latitudes the effective resolution is worse than this as the satellites view is less orthogonal with relation to the surface of the Earth. A parallax-corrected version of the data was used to reduce the distortion in the data. All missing values were replaced with zero and the data was scaled between zero and one.
The method developed in the study was benchmarked against the conventional methods. The first benchmark data is the output from the control member from MetCoOp Ensemble Prediction System (MEPS). MEPS is using Harmonie-AROME cycle 43h2.2 (Bengtsson et al., 2017), a non-hydrostatic, convection-permitting numerical weather prediction system run in cooperation by the meteorological institutes of Estonia, Finland, Norway and Sweden (Müller et al., 2017). MEPS forecast is produced every three hours and it provides an hourly forecast up to 66 hours, with a horizontal resolution of 2.5km. The cloud fraction is derived for each of the 65 vertical levels of the model using a statistical cloud scheme, and the total fraction is computed assuming a random overlap between successive layers. For verification the analysis field and the first five hours of the forecast were used. The second benchmark data was EXIM, an optical flow prediction system which is a part of the NWCSAF suite of products and (in FMI’s configuration) provides a 60-minute forecast in 15-minute prediction intervals (ZAMG, 2021). The horizontal resolution of EXIM is the same as the ground truth data, 3km. FMI archives contain EXIM data only for April and June 2024. The third benchmark data was Eulerian persistence: using the current cloudiness data as a (static) prediction field for all future predictions, which especially for the first intra-hour forecasts can be hard to beat.
@@ 2.2 Neural network
The CloudCast architecture was inspired by the U-Net model originally developed for image segmentation (Ronneberger et al., 2015). U-Net is an encoder-decoder convolutional neural network named after its symmetric U-shaped structure, which features a contracting path to capture context and an expanding path for localization. It employs skip connections, allowing input data to be directly transferred from the encoder to the decoder. Although initially developed for medical imaging, U-Net has proven to be a simple yet effective network, successfully applied to various domains. The advantages of U-Net over alternative convolutional networks include: 1) fast training due to its end-to-end structure and context-based learning, 2) skip connections that enable the reuse of low-level features, improving accuracy without requiring large datasets, and 3) the ability to capture both global context (via the encoder) and fine pixel-level details (via the decoder) (Siddique et al., 2021). U-Net has been widely adopted for many spatio-temporal and environmental forecasting tasks, often outperforming models with explicit temporal encoding (such as LSTMs or GRUs), as discussed in Section 1. This motivated the choice of U-Net, as it effectively captures spatial patterns while leveraging temporal context through multi-scale feature extraction. Furthermore, the relative simplicity of the U-Net makes it easier and faster to train compared to, for example, Generative Adversarial Networks (GANs) or diffusion-based networks. Given that the amount of satellite-based ground truth data available is limited to past six years, it is crucial to consider whether sufficient data exists to effectively train the network. Convolutional networks, which benefit from the strong inductive bias introduced by their convolutional layers, generally require less data than attention-based networks such as Vision Transformers (Cao et al., 2022). The U-Net architecture was implemented in CloudCast using the TensorFlow framework (Abadi et al., 2016).
@@
The network architecture is shown in Figure 3. The input comprises a sequence of effective cloudiness images, along with lead time and date/time information. As the data passes through the convolutional and max-pooling layers in the encoder, its resolution is gradually reduced, allowing for the extraction of increasingly abstract, high-level features. The encoder consists of four consecutive stages, during which the spatial dimensions are reduced to 32x32 pixels while the number of channels increases to 1024. Each stage includes convolutional layers followed by an activation function and a max-pooling layer to reduce spatial dimensions. The backbone of the network—situated between the encoder and decoder—also comprises convolutional layers followed by an activation function. The decoder mirrors the encoder’s structure but replaces convolutional layers with Conv2DTranspose layers to upsample the dense latent representation gradually back to the original spatial resolution. Additionally, the input data is concatenated directly with the decoder path to ensure the preservation of high-level features in the output. The final layer applies a sigmoid activation function to constrain the output values between zero and one, yielding a pixel-wise prediction of cloud cover for the specified lead time. The model comprises 31 million trainable parameters.
The approach of Sønderby et al. (2020) was adopted, where the forecast lead time is concatenated with the input data sequence. This method allows each lead time to be forecasted independently, enabling parallelization of the entire process and ensuring that each prediction is conditioned solely on ground truth data. However, using lead time as an additional input feature has drawbacks: the maximum lead time becomes an input parameter that must be determined prior to training the model, and the model’s ability to predict further into the future is constrained, as it does not observe the evolution of the weather system. While an autoregressive model can theoretically generate infinitely long predictions, the temporal range for this model is intentionally limited. This is acceptable as the focus of this study is exclusively on the nowcasting range of up to five hours.
@@ 2.3 Training details
To determine the optimal loss function and set of input features, systematic sensitivity experiments were conducted. These experiments tested different numbers of sequential input images—1, 2, 4, and 6—representing temporal histories of 0, 15, 45, and 75 minutes, respectively. Additionally, various static datasets were incorporated, including topography, ground type, date and time, and sun elevation angle. The strong bimodal distribution of the cloud cover data made selecting the right loss function particularly crucial: a total of six different loss functions were tested.
@@
Some of the widely used loss functions in machine learning, such as mean absolute error (MAE) and mean squared error (MSE), may not perform well with cloud cover, as they tend to minimize validation error by consistently predicting 50% cloud cover. They still serve as a very strong base line for the other loss functions. Binary cross-entropy (BCE) measures the difference between predicted and actual probability distributions and is typically used in binary classification tasks to predict the probability that an input belongs to one of two categories. In the case of CloudCast, the prediction is a two-dimensional array of cloud fraction values between zero and one. Notably, the distribution of these values is heavily concentrated at zero and one (see Figure 2). Because the cloud fraction values predominantly resemble a binary distribution, BCE can be used as a loss function, despite the non-categorical nature of the cloud cover data. Jungersen et al. (2022) found in their study BCE+MAE, a combination of BCE and MAE, to outperform plain BCE. The log-cosh function computes the logarithm of the hyperbolic cosine of the prediction error. It behaves like MSE for small error but is robust against outliers like MAE, making it suitable for time series prediction where outliers are present (Saleh and Saleh, 2022). Lastly, the Structural Similarity Index Measure (SSIM) assesses the similarity between two images based on structure, brightness, and contrast. It computes scores for multiple patches using a sliding window mechanism and aggregates these to provide a single global value for the image pair (Wang et al., 2004). As a loss function, SSIM evaluates structural similarity in images, emphasizing larger-scale spatial patterns rather than pixel-wise errors like MAE.
@@
The models trained in the sensitivity experiments are summarised in Table 1. Date and time were converted into cyclic variables using sine and cosine transformation. The terrain type was consolidated from 30 different categorical values into 6 major types: forest, urban, bare, permanent snow, water, agriculture, and other. The sun elevation angle was precomputed for each grid point and each hour of the year using an internal formula from FMI. Model names are abbreviated using convention loss function - number of input frames - additional features. Due to limited computation budget, the models were trained in a horizontal resolution of 20km.
Table 1: Models included in the sensitivity experiments. A total of 13 different models were trained. Model names are abbreviated using convention loss function-number of input frames-additional features, where the additional features are nd: no additional features, dt: date and time, terrain: terrain type, topo: topography and sun: sun elevation angle.
@@
The best model was selected based on the results of the sensitivity experiments and retrained using a 5km horizontal resolution (corresponding to 512x512 image size) over the same time period. The dataset size was 166GB. All the model hyperparameters were kept same, except batch size, which was reduced to 4. The training time for this model was 6 days.